# Fast constrained sampling in pre-trained diffusion models

### v2

@@ 3.1 Approximate Newton steps
10:     $\mathbf{\epsilon}_{t}=\frac{1}{\sqrt{1-\alpha_{t-s}}}{\bm{x}}_{t}-\frac{\sqrt{\alpha_{t-s}}}{\sqrt{1-\alpha_{t-s}}}\hat{{\bm{x}}}_{0}({\bm{x}}_{t})$
11:     ${\bm{x}}_{t-s}=\sqrt{\alpha_{t-s}}\hat{{\bm{x}}}_{0}({\bm{x}}_{t})+\sqrt{1-\alpha_{t-s}}\mathbf{\epsilon}_{t}+\beta_{t-s}{\bm{z}}_{t}$
12:  end for
13:  Return: ${\bm{x}}_{0}$
3.2 Symmetry of the Jacobian and difference in updates
Although, theoretically, the Jacobian of the denoiser should be symmetric (Eq. 14), the trained diffusion model does not exactly match the real score and can yield non-symmetric Jacobians. In Figure 2, we verify this claim with a simple experiment; we select a random image from the ImageNet (Deng et al., 2009) validation set and employ the Stable Diffusion 1.5 model (Podell et al., 2024) to denoise at three different noise levels $t=\{900,700,400\}$. We encode the image using the VAE encoder, scale it and add appropriate noise to get the intermediate diffusion latent for each timestep. We then give the noisy image to the denoiser network and compute the gradients $\partial\hat{{\bm{x}}}_{0}^{k,l}/\partial{\bm{x}}_{t}^{i,j}$ and $\partial\hat{{\bm{x}}}_{0}^{i,j}/\partial{\bm{x}}_{t}^{k,l}$ for randomly chosen pixels $(i,j),(k,l)$ using backpropagation. When we plot the gradients and compute the correlation coefficient $r$ we observe that the values deviate from $y=x$, which would indicate a symmetric Jacobian.
Having shown that the Jacobian is not symmetric, we expect to find differences between the gradient updates of gradient descent (Eq. 3) and our proposed update (Eq. 10). To highlight these differences, we set up a toy experiment where we update an identical initial image with the two different directions, under the same condition. The experiment is showcased in Figure 3 and again utilizes the Stable Diffusion 1.5 model. We create a synthetic black-and-white grid image (Figure 3 (a)), which we blur, deterministically encode, scale and noise to get a diffusion latent at $t=800$. We then set up a constraint where we add a red square to the center of the original image (Figure 3 (b)). The brown region in the constraint indicates masked-out values, or equivalently, where we will not be applying a constraint. Instead of decoding the image and applying the constraint in image space, we also encode the constraint image and apply it directly in the Stable Diffusion latent space.
We run 5 gradient updates with the same learning rate of 1 for each and demonstrate the final predicted $\hat{{\bm{x}}}_{0}$ for the ${\bm{J}}^{T}{\bm{e}}$ update of (3) (Figure 3 (c)) and the proposed ${\bm{J}}{\bm{e}}$ (10) (Figure 3 (d)). The final result shows how the model intends to change the entire image when asked to add a red square in the middle. We find that the resulting images differ substantially, with the proposed direction producing a more coherent image that tries to copy the newly introduced texture to the correct locations, i.e. the intersections of the lines. Although this is an empirical observation, we hypothesize that the two different directions can have vastly different effects on the image. In the Appendix A.2, we further visualize differences between ${\bm{J}}^{T}{\bm{e}}$ and ${\bm{J}}{\bm{e}}$, and in C.5, we repeat the previous experiment over multiple seeds.
4 Experiments
4.1 Linear Constraints
Figure 4: Comparison between our method and existing algorithms on free-form inpainting and 8×8$\times$8 × super-resolution. We directly use the images and results from (Chung et al., 2024) since there is no code available to replicate their method.
Inpaint (Free-form) SR (×$\times 8$)
Method PSNR ↑ LPIPS ↓ FID ↓ PSNR ↑ LPIPS ↓ FID ↓ Time
P2L (Chung et al., 2024) 21.99 0.229 32.82 23.38 0.386 51.81 30min*
LDPS 21.54 0.332 46.72 23.21 0.475 61.09 6min
PSLD (Rout et al., 2023) 20.92 0.251 40.57 23.17 0.471 60.81 8min
Ours 21.73 0.258 19.39 24.26 0.455 60.99 15s/1min
Table 1: Quantitative evaluation (PSNR, LPIPS, FID) on free-form inpainting and 8×8$\times$8 × superresolution on ImageNet. Our algorithm is competitive with the results of P2L which requires at least 30×30$\times$30 × more time for a single sample.
Figure 5: Qualitative evaluation of large area (box) inpainting on ImageNet. Our method provides results close to the fine-tuned inpainting model while requiring a fraction of the time to run per image.
We first verify our algorithm by generating images under linear constraints, which has been the main application of many previous algorithms (Chung et al., 2023; Rout et al., 2023; Chung et al., 2024). We follow the evaluation setting of Saharia et al. (2022a) and test our method on ImageNet (Deng et al., 2009), using the first 1000 images from the 10k validation set of the ctest10k split. As our evaluation metrics, we measure the PSNR and LPIPS (Zhang et al., 2018) between the generated and real images, as well as the FID (Heusel et al., 2017) between the real and generated sets. We use Stable Diffusion 1.4, which is pre-trained on the LAION (Schuhmann et al., 2022) text-image pair dataset. All experiments were run on NVIDIA RTX A5000 24GB GPUs.
Free-form Inpainting & Super-resolution
We follow the previously described experiment settings of free-form inpainting and 8×8$\times$8 × super-resolution. For free-form inpainting, the masks are randomly sampled and mask out 10-20% of the image pixels, as done by Saharia et al. (2022a). For inpainting with our method, we opt to directly operate on the Stable Diffusion latent given that Stable Diffusion VAE mostly compresses information locally. We apply the masking in pixel space, encode the masked image and inpaint with the unmasked VAE latents. We also apply a $3\times 3$ dilation kernel on the pixel mask before downsampling, masking out some extra pixels along the edge that we find the VAE fails to encode. For inpainting, we set the number of optimization steps $K=5$ over which we linearly decrease the learning rate $\lambda$ from 0.5 to 0.1. For super-resolution, we use $K=10$ and a constant $\lambda=0.1$. We perform only 20 DDIM (Song et al., 2020) steps for all experiments.
Regarding super-resolution, we cannot apply the constraint in the VAE latent space as image pixel downsampling does not correspond to downsampling VAE latents. To compute the error direction ${\bm{e}}$ we backpropagate the pixel-level constraint cost $({\bm{A}}\mathcal{D}({\bm{x}}_{t})-{\bm{y}})^{T}({\bm{A}}D({\bm{x}}_{t})-{\bm{y}})$ through the decoder network and utilize the resulting gradient w.r.t. the latent as the error direction. We note that by setting ${\bm{e}}=-\partial C(\mathcal{D}({\bm{x}}_{t}))/\partial{\bm{x}}_{t}$ we only require backpropagation through the decoder model, which is significantly less expensive than backpropagating through the denoiser network, as done by previous works. For both degradations, we also include additive white Gaussian noise with $\sigma_{\bm{y}}=0.05$.
We present the results on free-form inpainting and 8×8$\times$8 × super-resolution (SR) in Table 1. In inpainting, we find that our method generates better-aligned parts for the missing image regions, reflected in the significant improvement in FID, while obtaining similar PSNR and LPIPS scores to state-of-the-art methods. For super-resolution, although the improvements are smaller, we still find that we can compete with existing methods in both the quality and faithfulness of the generated images.
The main advantage we showcase is inference time; our approach achieves similar or better results to every other existing method while only requiring a fraction of the computation. When we do not perform backpropagation through the VAE decoder (inpainting), our method requires only 15 seconds. In super-resolution, where we perform more optimization steps and backpropagate, the time increases to 1 minute. In contrast, P2L (Chung et al., 2024), the best-performing method, requires 30 minutes per single image since it relies on additional backpropagation steps through the denoiser to infer the text condition. Since there is no publicly available implementation for P2L we report the results directly from the paper and estimate the inference time as 5×5$\times$5 × that of LDPS (Chung et al., 2023), which the authors mentioned as a reasonable expectation as it linearly scales with the number of text prompt optimization iterations.
Box Inpainting
In the previous experiments, we observed that our method performed better in inpainting, which required the model to infer more missing information in the given image than in super-resolution. We further push the model by asking it to inpaint a box that covers 25% of the input image. In the comparisons we also include FreeDoM (Yu et al., 2023), which, although not previously shown on inpainting, claims to be a fast training-free inference approach for any condition. We also include the fine-tuned SD-Inpaint model (Rombach et al., 2022), which required 500k additional training steps.
We present the results on box inpainting in Table 2 and Figure 5. Our approach outperforms all existing training-free methods in all metrics and is the closest to the fine-tuned SD-Inpaint model, which we consider an upper limit. Qualitatively, we find that LDPS and PSLD, which perform a lot of denoising steps, generate blurry parts that align with the average appearance over the rest of the image. FreeDoM, the faster baseline, although generating non-blurry parts, frequently fails to maintain consistency with the rest of the image. Our method achieves both high quality and consistency in a shorter time as all previous methods backpropagate through the denoiser weights. In Appendix C.1 we ablate the steps $K$ and learning rate $\lambda$ hyperparameters.
Inpaint (Box)
Method PSNR ↑ LPIPS ↓ FID ↓ Time
LDPS 17.52 0.42 76.32 6min
PSLD 17.30 0.38 74.02 8min
FreeDoM 16.18 0.42 55.68 1min
(Yu et al., 2023)
Ours 18.30 0.30 42.01 15s
SD-Inpaint 19.05 0.28 32.93 4s
Table 2: Quantitative evaluation (PSNR, LPIPS, FID) on large area (box) inpainting.
4.2 Style-guided generation
Our algorithm can be applied to non-linear differentiable constraints using backpropagation to approximate ${\bm{e}}$. Computing this gradient direction does not require backpropagation through the denoiser but through the separate, more lightweight ’critic’ network $f$. We demonstrate this with the non-linear constraint of style-guided generation.
The goal is to generate an image that simultaneously follows the style of a reference image ${\bm{x}}_{ref}$ and a given text prompt. Following previous works (Yu et al., 2023; He et al., 2024) which also perform style-guided generation, we do this by matching the statistics of CLIP (Radford et al., 2021) features between the reference and the generated images. We minimize the Frobenius norm between the Gram matrices (Gatys et al., 2016) of the 2nd layer CLIP features, making the cost to be minimized
$\|\text{Gram}(\text{CLIP}_{2}({\bm{x}}_{ref}))-\text{Gram}(\text{CLIP}_{2}(\hat{{\bm{x}}}_{0}))\|_{F}^{2}.$ (15)
Adhering to the evaluation of Yu et al (2023), we use 1000 random pairs of reference style images from WikiArt (Tan et al., 2019) and prompts from PartiPromtps (Yu et al., 2022). We measure the CLIP similarity between the generated images and the text prompts to evaluate the faithfulness to the text condition, and the final difference between the Gram matrices to evaluate the faithfulness to the style reference (Style Score). We use the CLIP ViT-B/16 model for guiding the style of the image and evaluating. We also repeat the experiment using the OpenCLIP ViT-B/32 model (Cherti et al., 2023) for guidance. We perform $K=5$ gradient updates for every denoising step, using a linearly decreasing learning rate $\lambda$ from 0.5 to 0.1 and classifier-free guidance (Ho & Salimans, 2022) $w=2$ and $w=5$ for the denoiser.
The results are presented in Table 3 and Figure 6. We find that our method achieves the overall best style score. Our approach is the best at minimizing the constraint which is also the the evaluation metric. Although in general, the lower the style score the more difficult it is to maintain high CLIP similarity with the prompt, we observe that our method still balances between the style and text, especially when we increase the guidance weight. We also used an OpenCLIP model to guide the style. With a CLIP model still as the evaluator, we achieve better style scores than the baselines that optimized directly with CLIP. Our proposed update reduces the target cost better without generating adversarial artifacts that trick the CLIP evaluation.
Regarding inference time, our method is the fastest out of all baselines, excluding the non-style-guided DDIM. The baseline of MPGD (He et al., 2024), which is comparable in inference speed, does not backpropagate the style loss through the denoiser network, but only modifies the $\hat{{\bm{x}}}_{0}$ estimation at every denoising step. As we demonstrate in the supplementary, MPGD completely fails to propagate the constraint to distant pixels when the conditioning affects a specific region in the image. This makes it unusable in some constrained sampling settings, such as box inpainting.
Figure 6: Style-guided generation. We show the results of guiding Stable Diffusion with a CLIP and an OpenCLIP model, using classifier-free guidance $w=5$. The images generated by our algorithm follow the reference style better than previous algorithms.
Method Style Score* ↓ CLIP ↑ Time
DDIM 761.0 31.61 9s
FreeDoM 498.8 30.14 80s
MPGD (He et al., 2024) 441.0 26.61 50s
Ours (w=2) 368.37 23.95 45s
Ours (w=5) 310.96 24.57 45s
Ours (OpenCLIP, w=5) 434.45 25.94 45s
Table 3: Quantitative evaluation of style generation. *The Style Score is what the gradient is optimizing.
4.3 Mask-guided generation
As an additional non-linear constraint, we guide the Stable Diffusion model with a separately trained face segmentation network. We employ an off-the-shelf model from huggingface and set the constraint $C$ to be the KL divergence between the per-pixel segmentation classes predicted for a reference image and the generated image. Using 100 images from the FFHQ (Karras et al., 2019) validation set, we run both our method and the MPGD (He et al., 2024) baseline. We chose MPGD because as the fastest baseline that works very well with ’dense’ constraints, i.e. constraints that are applied to all pixels.
The results in Table 4 show that with a similar compute budget, our method achieves both faithfulness (mIoU between generated and reference images) and image quality (CLIP-FID). We used CLIP-FID because it performs better than Inception-FID on a small set of images (Kynkäänniemi et al., 2023). MPGD with a high weight ($\rho$) ’burns in’ the segmentation mask, leading to non-realistic images, whereas a lower weight does not produce images faithful to the mask. We provide qualitative examples of the generated images and the masks used in the Appendix C.5.
Method mIoU ↑ CLIP-FID ↓ Time
DDIM 0.09 48.78 3s
MPGD ($\rho=1$) 0.47 77.11 1min
MPGD ($\rho=0.5$) 0.36 56.38 1min
Ours 0.42 59.79 1min
5 Conclusion
In this work, we presented a new algorithm for fast inference in pre-trained diffusion models under constraints. Our approach exploits the relationship between the noisy and clean images during the denoising process to perform a gradient update that requires no expensive backpropagation operations through the diffusion model. Our method produces results comparable to the state-of-the-art with a fraction of the inference time. Our approach offers a practical method to sample from large pre-trained generative image models under any condition and has the potential to enable training-free downstream applications that rely on utilizing a strong image prior.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. We acknowledge that there are many potential societal consequences of our work. Beyond the positives, we highlight the ability to edit images with the intent to deceive and mislead as requiring careful consideration.
Acknowledgments
Part of this work was done during an internship at Microsoft Research, Redmond. This research was also partially supported by NSF grants IIS-2123920, IIS-2212046.
Appendix A Analysis of the proposed algorithm
A.1 Convergence
We resort to inexact Newton methods (Dembo et al., 1982) to analyze the convergence and validity of the proposed approach. The original Newton’s method aims to solve the system
${\bm{J}}{\bm{e}}_{t}={\bm{e}}.$ (16)
In cases where computing the inverse of the Jacobian, as is the case of the diffusion denoiser Jacobian, inexact Newton methods propose first finding an approximation ${\bm{e}}_{t}^{*}$ to the solution of Eq (16) and then performing the Newton step using the approximate solution and reiterating until convergence. For inexact Newton methods to converge to the correct solution, the residual must be strictly reduced at every step
If we substitute the proposed update ${\bm{e}}_{t}^{*}=\lambda{\bm{J}}{\bm{e}}$ (Eq. 10) we get the residual
${\bm{r}}={\bm{e}}-\lambda{\bm{J}}^{2}{\bm{e}}=({\bm{I}}-\lambda{\bm{J}}^{2}){\bm{e}}.$ (18)
For this residual to be strictly reduced at every iteration of our algorithm, we need
Therefore, we need to show that the spectral norm of the matrix ${\bm{I}}-\lambda{\bm{J}}^{2}$ is smaller than 1 for a correct choice of learning rate $\lambda$.
If ${\bm{J}}$ was diagonalizable (or more generally normal) then we could directly estimate $\lVert{\bm{I}}-\lambda{\bm{J}}^{2}\rVert_{2}$ using the largest eigenvalue of ${\bm{J}}$, following the spectral theorem. While we cannot directly assume that ${\bm{J}}$ is normal, we know that ${\bm{J}}$ should be ’almost’ symmetric, as is the optimal denoiser solution shown in Eq. 14. This means that we could express ${\bm{J}}={\bm{S}}+{\bm{K}}$, where ${\bm{S}}=({\bm{J}}+{\bm{J}}^{T})/2$ is the symmetric component and ${\bm{K}}=({\bm{J}}-{\bm{J}}^{T})/2$ the skew-symmetric component, and expect $\lVert{\bm{K}}\rVert\ll\lVert{\bm{S}}\rVert$. In that case, we can approximate $\lVert{\bm{I}}-\lambda{\bm{J}}^{2}\rVert_{2}$ as
where $\mu_{i}$ are the eigenvalues of ${\bm{S}}$.
To get an estimate of the magnitudes of the largest eigenvalues of ${\bm{S}}$, which are real, and of the eigenvalues of ${\bm{K}}$, which are imaginary, we use the Arnoldi iteration (Arnoldi, 1951). For the Arnoldi iteration we only require access to the matrix-vector products ${\bm{J}}{\bm{v}}$ and ${\bm{J}}^{T}{\bm{v}}$, which can be computed using the approximation of Eq. (12) and backpropagation respectively.
If we approximate the spectral norm of Eq.( 19) using the symmetric component of the Jacobian we see that for a correct choice of $\lambda$ we can ensure that $\lvert 1-\lambda\max_{i}\mu_{i}\rvert<1$. Now we revisit the example of Fig. 1 where we inpaint half of the image. Instead of using the empirical learning rate used in the paper (linearly decreasing and gradient normalized by the $\infty$-norm) we compute the largest eigenvalue of $({\bm{J}}+{\bm{J}}^{T})/2$, $\mu$, and select different $\lambda$ so that we satisfy or violate the bound provided by Eq. (20).
A.2 Difference between ${\bm{J}}$ and ${\bm{J}}^{T}$
In the main text, we referred to the difference between ${\bm{J}}$ and ${\bm{J}}^{T}$ as a motivating factor for our approach and measured that difference by comparing elements $(i,j)$ and $(j,i)$ of the Jacobian matrix (Figure 2). In Figure 8 we further visualize the difference between using the Jacobian and its transpose by numerically computing how the Jacobian and the Jacobian transpose of the entire image would look for a single pixel, i.e. which parts of the image are affected by a change in a single pixel.
Contrary to the expectations, we find a significant difference between our proposed direction and backpropagation, with our approach having a better effect on retaining shapes and symmetry across the image. We see that in many cases, the model is trying to change ’corresponding’ parts of the image together, i.e. change both eyes or ears simultaneously, giving us a better understanding of the knowledge that the model has acquired regarding the image space in general.
Appendix B Proposed algorithm as modification of DDIM
To apply the proposed method we modified the DDIM sampling algorithm (Song et al., 2020). We provide a side-by-side comparison to show the difference between the original DDIM and the proposed algorithm. Extending our algorithm to other sampling methods should be intuitive by alternating between gradient updates from our algorithm and the diffusion updates computed with the diffusion sampling algorithm used.
Algorithm 2 DDIM sampling (Song et al., 2020) using a pre-trained diffusion model.
1:  Input: Pre-trained diffusion model $\hat{{\bm{x}}}_{0}({\bm{x}}_{t})$, diffusion step size $s$ and schedule parameters $\alpha_{i},\beta_{i}$
2:  ${\bm{x}}_{T}\sim N(\textbf{0},{\bm{I}})$
3:  for $t\in\{T,T-s,T-2s,\dots,s\}$ do
4:     ${\bm{z}}_{t}\sim N(\textbf{0},{\bm{I}})$
5:     $\mathbf{\epsilon}_{t}=\frac{1}{\sqrt{1-\alpha_{t-s}}}{\bm{x}}_{t}-\frac{\sqrt{\alpha_{t-s}}}{\sqrt{1-\alpha_{t-s}}}\hat{{\bm{x}}}_{0}({\bm{x}}_{t})$
6:     ${\bm{x}}_{t-s}=\sqrt{\alpha_{t-s}}\hat{{\bm{x}}}_{0}({\bm{x}}_{t})+\sqrt{1-\alpha_{t-s}}\mathbf{\epsilon}_{t}+\beta_{t-s}{\bm{z}}_{t}$
7:  end for
8:  Return: ${\bm{x}}_{0}$
Algorithm 3 The proposed algorithm for constrained sampling using a pre-trained diffusion model.
1:  Input: Pre-trained diffusion model $\hat{{\bm{x}}}_{0}({\bm{x}}_{t})$, constraint $C({\bm{x}}_{0},{\bm{y}})=\lVert f(\hat{{\bm{x}}}_{0}({\bm{x}}_{t})-y\rVert_{2}^{2}$, condition ${\bm{y}}$, gradient step size $\delta$, iterations $K$, learning rate $\lambda$, diffusion step size $s$ and schedule parameters $\alpha_{i},\beta_{i}$
2:  ${\bm{x}}_{T}\sim N(\textbf{0},{\bm{I}})$
3:  for $t\in\{T,T-s,T-2s,\dots,s\}$ do
4:     for $i\in\{1,2,\dots,K\}$ do
5:        ${\bm{e}}={\bm{J}}_{f}^{-1}(f(\hat{{\bm{x}}}_{0}({\bm{x}}_{t}))-{\bm{y}})$Task-dependent}
7:        ${\bm{x}}_{t}={\bm{x}}_{t}+\lambda{\bm{e}}_{t}$
8:     end for
9:     ${\bm{z}}_{t}\sim N(\textbf{0},{\bm{I}})$
10:     $\mathbf{\epsilon}_{t}=\frac{1}{\sqrt{1-\alpha_{t-s}}}{\bm{x}}_{t}-\frac{\sqrt{\alpha_{t-s}}}{\sqrt{1-\alpha_{t-s}}}\hat{{\bm{x}}}_{0}({\bm{x}}_{t})$
11:     ${\bm{x}}_{t-s}=\sqrt{\alpha_{t-s}}\hat{{\bm{x}}}_{0}({\bm{x}}_{t})+\sqrt{1-\alpha_{t-s}}\mathbf{\epsilon}_{t}+\beta_{t-s}{\bm{z}}_{t}$
12:  end for
13:  Return: ${\bm{x}}_{0}$
Appendix C Additional Results
C.1 Number of steps and learning rate ablation
We repeat the ImageNet box inpainting experiments with fewer optimization iterations and a higher learning rate. We show qualitative results in Figure 9, where we observe that running fewer optimization steps gives blurrier results, which is expected as the known regions of the image also seem to not have converged to the given values. Using a higher learning rate leads to the model sometimes ’overshooting’ by inpainting the missing regions with realistic-looking parts that do not necessarily fit the rest of the image. The quantitative results are presented in Table 5.
Method PSNR ↑ LPIPS ↓ FID ↓
Ours (n=5, lr=0.5) 18.30 0.30 42.01
Ours (n=2, lr=0.5) 18.01 0.39 68.75
Ours (n=5, lr=1.0) 17.48 0.32 47.20
Table 5: Ablation of the number of steps and learning rate on box inpainting. When using fewer steps, the algorithm does not sufficiently converge. When using a very high learning rate, the approach seems to overshoot to non-realistic inpainted parts.
@@ C.2 Non-differentiable constraints
As a toy example, we generate images with pixel values quantized to be either ’on’ or ’off’ (-1 or 1). The constraint first measures whether a pixel value is positive or negative and then sets the error direction ${\bm{e}}$ to $|1-x|$ or $|-1-x|$ for every pixel accordingly. This is a non-differentiable constraint for which we can easily compute a local gradient that reduces the cost $C$. Using the prompt ’a photo of a cat’, we generate quantized images as shown in Figure 10.
@@
C.3 Inference visualizations
In Figures 11, 12, and 13, we visualize the intermediate steps of the proposed algorithm for the inpainting, super-resolution and style-guided generation tasks respectively. Our method quickly converges to a plausible image and then further refines it to better satisfy the constraint over the diffusion timesteps. For style-guided generation, we see that the structure of the image is defined in the first few initial steps before the specific style provided is applied.

### v3

@@ A.6 DDIM and other sampling methods
10:  ${\bm{x}}_{t+1}=\text{Solver}(\hat{{\bm{x}}}_{0}({\bm{x}}_{t}),{\bm{z}}_{t},\alpha_{t},t+1)$
11:end for
12:Return: ${\bm{x}}_{N}$
Appendix B Additional results
B.1 Mask-guided generation
For the mask-guided generation experiment we utilized a pre-trained face segmentation model from huggingface https://huggingface.co/jonathandinu/face-parsing. In Figure 13 we provide qualitative results of our segmentation mask-guided generation experiment described in the main text. The quantitative results are provided in Table 5 in the main text.
B.2 Number of steps and learning rate ablation
We repeat the ImageNet box inpainting experiments with fewer optimization iterations and a higher learning rate. We show qualitative results in Figure 14, where we observe that running fewer optimization steps gives blurrier results, which is expected as the known regions of the image also seem to not have converged to the given values. Using a higher learning rate leads to the model sometimes ’overshooting’ by inpainting the missing regions with realistic-looking parts that do not necessarily fit the rest of the image. The quantitative results are presented in Table 4 in the main text.
@@ B.3 Non-differentiable constraints
As a toy example, we generate images with pixel values quantized to be either ’on’ or ’off’ (-1 or 1). The constraint first measures whether a pixel value is positive or negative and then sets the error direction $\bm{e}$} to |1x||1-x| or |1x||-1-x| for every pixel accordingly. This is a non-differentiable constraint for which we can easily compute a local gradient that reduces the cost $C$. Using the prompt ’a photo of a cat’, we generate quantized images as shown in Figure 15.
@@
B.4 Time-distilled models
We test whether our method works on distilled models such as rectified flows [20] and consistency models [35]. These techniques reduce the number of inference steps by distilling from a base diffusion model. Our method does not explicitly depend on the number of inference steps used, the type of model or the noise schedule; the only requirement is having a way to estimate the final clean image from the current step, which both rectified flows and consistency models admit. Thus, applying to rectified flows and consistency models is intuitive.
We employ our method to inpaint images using the 2-rectified flow model distilled from Stable Diffusion, InstaFlow [21], using 5 inference steps. We observe that although our algorithm still works with a rectified flow and just 5 inference steps, we do not consistently get high-quality samples as we did with Stable Diffusion when using the same hypeparameters and it requires more optimization steps. Ultimately, we utilize line search to find the $\lambda$ for every gradient update we perform. This increases inference time further, mitigating the gains from using a time-distilled model.
We hypothesize that using rectified flows (or any distilled model with fewer steps) may be more challenging since the initial noise dictates most of the content in the final image. Therefore, the first optimization steps we perform must get sufficiently close to the correct solution.
When using a diffusion model, we can get away with imperfect optimization steps as there is more room for ’fixing’ the image in later timesteps. We show examples of the rectified flow inpainting in Figure 16. We also refer the reviewer to Figures 17,18, where we show the intermediate generation steps for Stable Diffusion.
Figure 16: Using a rectified flow model [21] that generates images in just 5 steps requires more careful optimization steps, as there is less room for errors during generation.
B.5 Inference visualizations
In Figures 17, 18, and 19, we visualize the intermediate steps of the proposed algorithm for the inpainting, super-resolution and style-guided generation tasks respectively. Our method quickly converges to a plausible image and then further refines it to better satisfy the constraint over the diffusion timesteps. For style-guided generation, we see that the structure of the image is defined in the first few initial steps before the specific style provided is applied.